OpenStudy (anonymous):

Indicate which property is illustrated in Step 3. Step 1 10 - 6x - 10 + 7x = (10 - 6x) + (-10 + 7x) Step 2 = (10 - 6x) + (7x - 10) Step 3 = 10 + (-6x + 7x) - 10 Step 4 = 10 + (-6 + 7)x - 10 Step 5 = 10 + x - 10 Step 6 = x + 10 - 10 Step 7 = x + 0 Step 8 = x commutative associative distributive identity

OpenStudy (anonymous):

@e.mccormick @campbell_st @mathmath333 @aaronq

OpenStudy (anonymous):

I think it is associative.

OpenStudy (anonymous):

Indicate which property is illustrated in Step 1. Step 1 8 - 7x - 8 + 8x = (8 - 7x) + (-8 + 8x) Step 2 = (8 - 7x) + (8x - 8) Step 3 = 8 + (-7x + 8x) - 8 Step 4 = 8 + (-7 + 8)x - 8 Step 5 = 8 + x - 8 Step 6 = x + 8 - 8 Step 7 = x + 0 Step 8 = x distributive identity associative commutative

OpenStudy (anonymous):

A?

OpenStudy (anonymous):

@alyisnumber1

OpenStudy (anonymous):

@campbell_st

OpenStudy (campbell_st):

sorry its just a bit bizarre for me...

OpenStudy (anonymous):

So you don't know which it is?

OpenStudy (anonymous):

Commutative.